The town of Recycleville has 30,000 citizens.Each home in Recycleville generates 1.1 tons of solid waste per year and has an average of 2.4 residents per service customer.The town has no commercial businesses or schools.The local landfill has space for 100,000 tons and charges $10 per ton.Once the landfill is full,the waste will need to be sent to a remote landfill at a cost of $90 per ton.What is the 10-year average disposal cost for the town,assuming a rate of inflation of 5% per year?what will be the cost per home?
Answer #1

Let no. of citizen per home = 4
Total no. of home = 30000 / 4 = 7500 homes
Total wastage = 1.1 x 7500 = 8250 tons per year

1. Total wastage in 10 years @5% inflation = 8250 x (1.05)^ = 13438.3 tons

Total cost of land space = 13438.3 x 10 = $134383

Cost of transfer to remote landfill = 13438.3 x $90 = $1209447

Total cost = $134383 + $1209447 = $1343830

2. Cost per home = Total cost / no. of homes = $1343830 / 7500 = $179.2

Average cost per citizen = $1343830 / 30000 = $44.8 per citizen